solve each system using the substitution method or the addition method
3(2x+y)=4x+20
x-2y=3

Answer by mananth(16949)

3(2x+y)=4x+20
6x +3y =4x+20
2x+3y=20
x-2y=3

2 x + 3.00 y = 20 .............1
Total value
1 x -2 y = 3 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 3.00
4 x 6 y = 40
3 x -6 y = 9
Add the two equations
7 x = 49
/ 7
x = 7
plug value of x in (1)
2 x + 3 y = 20
14 + 3 y = 20
3 y = 20 -14
3 y = 6
y = 2
x= 7
y= 2
m.ananth@hotmail.ca